Jennifer bought jshares of Keester's Koverups for ndollars last year. Her stock rose in price and eventually would earn her 120% capital gain. She decided to sell half of her shares. Represent the capital gain earned on each of the shares that were sold algebraically.

a. 1.2n

b. 0.5n

C. 120j

d.0.5j


1
Expert's answer
2021-09-30T15:01:12-0400

Let Jennifer bought (x) shares of Keester's for (n) dollars. 

Half of these shares ("\frac{x}{2}" ) cost in present = "\frac{n}{2}" dollars .

Now , her stocks appreciate by 120%

So , 

Appreciated Price of (x) stocks = "(1 + 120\%) n"

"= 2.2 n"

Price she gets for selling half the shares "\frac{x}{2} = \frac{2.2n}{2} = 1.1n"

 

Capital Gains for selling half the shares

"=\frac{Value\space of\space half \space stocks\space after\space appreciation - Value\space of\space Half\space stocks\space initially}{Value \space of\space Half\space stocks\space initially}"

"=\frac{1.1n - 0.5n}{0.5n}\\

= 1.2n"